# Rethinking the Learning Paradigm for Facial Expression Recognition

### v2

@@ III-C Label Disambiguation
The label disambiguation with the transformer decoder includes the label embedding regularization module and the confidence revision module. An input image $x$ is fed into the pre-trained encoder to get visual tokens $\mathcal{F}^{HW\times d}$, where $HW$ denotes the number of visual tokens and $d$ denotes the feature dimension of the hidden state. We initialize the query embedding $\mathcal{Q}^{k\times d}$, where $k$ denotes the number of classes. In the fine-tuning stage, the transformer decoder updates $\mathcal{Q}^{k\times d}$ according to visual tokens, with $\mathcal{Q}^{k\times d}$ reflecting the correlation of category-related features. Inspired by [52, 53], we link the category-related features to the learnable query embedding $\mathcal{Q}^{k\times d}$ and predict the confidence sector $C_{i}$ of the sample by their correspondence. Specifically, $C_{i}$ is obtained by inputting the logits from the last layer of the transformer decoder to the sigmoid function. However, differently by [52, 53], (1) our method is designed for PLL and focuses on disambiguation by query to finally find the unique correct label with the highest confidence for each instance. (2) Besides querying the classification logits, we also query the label confidence. (3) To mitigate the problem of imbalance distribution in the FER dataset, we design a special regularization for label embedding.
III-C1 Label Embedding Regularization.
@@
It is worth noting that in addition to the annotation ambiguity, these FER datasets also commonly exhibit imbalance distribution. In the training stage, it makes the model more biased towards majority classes, resulting in an uneven distribution of the feature space, and degrading model performance by classifying minority classes as head classes. In our work, we draw inspiration from  [54] and aim to achieve a uniform distribution of the query embedding over a hypersphere to mitigate the negative effects of class imbalance. Specifically, we compute a uniform loss for query embedding to pull the distance between different classes on the hypersphere. We compute the optimal positions for different classes of query embedding $\left\{t_{i}\right\}_{i=1}^{K}$ before training and optimize it using the following (eq. 2):
$\mathcal{L}_{uniform}\left(\left\{t_{i}\right\}_{i=1}^{K}\right)=\frac{1}{K}\sum_{i=1}^{K}\log\sum_{j=1}^{K}e^{t_{i}^{T}\cdot t_{j}/\tau},$ (2)
@@
where $\tau>0$ is a temperature hyperparameter, and $K$ denotes the number of classes in FER dataset.
III-C2 Revision Confidence.
@@
As previously mentioned, the robust representation in the feature space enhances PLL’s ability to disambiguate in the label space. For samples with ambiguous labels in the dataset, we initially set the candidate label set $Y$ with a uniform distribution, represented by $C_{i}=\frac{1}{|Y_{i}|}Y_{i}$. During the update process, we continuously revise the confidence score $C_{i}$ for each batch using the following (eq. 3):
$\binom{id_{1}:v_{1}}{id_{2}:v_{2}}=Top_{k}[\varphi*C_{i-1}],k=2,$ (3)
@@
where $\varphi=sigmoid(\sigma_{logits})$ is the response of sigmoid layer, and $\sigma_{logits}$ is the output of the transformer decoder. After that, we compute the top-2 metrics of the $(C_{i-1}*\varphi)$ for updating confidence vector $C_{i}$ as the (eq. 4):
@@
$C_{i}=\left\{\begin{array}[]{lcr}C_{i-1}\;\;\quad\quad\quad\quad\quad\quad\Delta{v}\leq threshold\\ C_{i-1}[id_{1}]=1\quad\quad\quad else\end{array}\right.,$ (4)
@@
where $\Delta{v}=|v_{1}-v_{2}|$, the threshold $\in$ [0,1). When $\Delta{v}>$ threshold, $C_{i}$ is a one-hot vector, where only the 1-th entry is 1. We treat it as the level of uncertainty and represent it with the top-2 rank label in the candidate label set. algorithm 1 is the pseudo-code of our method, which would converge to a stable status along the training iterations. We specifically choose top-2, because we notice that the accuracy of most samples located in top-2 reaches 95%.
@@
0:  $Model\;f$: Pre-trained model $f$;
@@
  $query\;embedding$: query $\mathcal{Q}$ ;
@@
        calculate Uniform distribution for query by eq. 2);
        revise confidence vector $C_{i}$ by eq. 3) and eq. 4);
        calculate PLL loss by eq. 1)
        update Network parameter of $f$ by eq. 1);
@@
IV Experiments
In this section, we describe the FER datasets in section IV-A, implementation details in section IV-B, pre-training, and fine-tuning details in section IV-C, comparison with Fer methods in section IV-D, ablation studies in section IV-E and visualization results in section IV-F.
To evaluate our proposed framework, we compare it with a baseline and other related methods including fully-supervised learning: PSR [25], DDA [26], SCN [20], SCAN [27], DMUE [29], RUL [28], DACL [55], MA-Net [56], MVT [57], VTFF [58], EAC [59], and TransFER [60], LA-Net [61]; and partial label learning: RC [62], LW [38], PICO [42], and CRPLL [63].
IV-A Datasets.
RAF-DB [64] contains around 30K diverse facial images with single or compound expressions labeled by 40 trained annotators via crowdsourcing. In our experiments, we use 12271 images for training and 3068 images for testing. FERPlus [9] is an extended dataset of the standard emotion FER dataset with 28709 training, 3589 validation, and 3589 test images. It provides a set of new labels for each image labeled by 10 human annotators. According to previous work, we just use the train part and test part. AffectNet [11] is one of the most challenging FER datasets with manually labeled 440k images. As per prior research, we divide it into AffectNet-7 and AffectNet-8, with the latter having one more expression of contempt and 3667 training and 500 test images more compared to the former.
TABLE I: Comparison with the state-of-the-art results on the RAF-DB, FERPlus, AffectNet7, and AffectNet8 datasets. Accuracy (Acc.(%)) is reported on the test dataset.
Method Backbone RAF-DB FERPlus AffectNet7 AffectNet8
PSR [25] VGG-16 88.98 89.75 63.77 60.68
DDA [26] ResNet-18 86.90 - 62.34 -
SCN [20] ResNet-18 87.03 88.01 63.40 60.23
SCAN [27] ResNet-50 89.02 89.42 65.14 61.73
DMUE [29] ResNet-18 88.76 88.64 - 62.84
RUL [28] ResNet-18 88.98 88.75 61.43 -
DACL [55] ResNet-18 87.78 88.39 65.20 60.75
MA-Net [56] ResNet-18 88.40 - 64.53 60.96
MVT [57] DeIT-S/16 88.62 89.22 64.57 61.40
VTFF [58] ResNet-18+ViT-B/32 88.14 88.81 64.80 61.85
EAC [59] ResNet-18 89.99 89.64 65.32 -
TransFER [60] CNN+ViT 90.91 89.76 66.23 -
LA-Net [65] ResNet-50 91.56 67.60
RC [62] ResNet-18 88.92 88.53 64.87 61.63
LW [38] ResNet-18 88.70 88.10 64.53 61.46
PICO [42] ResNet-18 88.53 87.93 64.35 61.38
CRPLL [63] ResNet-18 88.47 88.01 64.32 61.37
Baseline [44] ViT-B/16 88.36 88.18 64.61 61.34
Ours ViT-B/16 92.05 91.11 67.11 64.12
IV-B Baseline and Experiment Setup.
We adopt ViT-B/16 [44] pre-trained on Imagenet-1K as our baseline. For fairness, we keep the same parameters as our fine-tuning part in all settings, which will be described later. In terms of SL methods, we adopt their optimal settings to make fair comparisons. For data preprocessing, we keep all the images with the size of $224\times 224$ and use MobileFacenet [66] to obtain aligned face regions for RAF-DB and FERPlus. For AffectNet, alignment was obtained via the landmarks provided in the data. In the pre-training part, we use four Tesla V100 GPUs to pre-train the ViT for 600 epochs with batch size 192 and learning rate 2e-4. We set the mask ratio of Maskfeat to 0.4. In the fine-tuning part, we use only two Tesla V100 GPUs. We utilize the AdamW optimizer, setting the batch size to 320 and the initial learning rate to 1e-4 with the weight decay of 5e-2. The pre-training and fine-tuning protocol that we use is based on [50].
IV-C Details for pre-training and fine-tuning
table II and table III summarize the pre-training configurations of different pre-trained methods. LABEL:fine-1 shows the fine-tuning configurations on the FER dataset. We use three FER datasets to pre-train the encoder part and adopt 2 transformer blocks in the decoder part.
TABLE II: Pre-training setting, Part I.
@@ IV-C Details for pre-training and fine-tuning
TABLE III: Pre-training setting, Part II.
@@
TABLE IV: Fine-tuning setting of Maskfeat on FER datasets.
@@
IV-D Comparison with the SOTA FER Methods
@@
We conduct extensive comparisons with the SOTA FER methods on RAF-DB, FERPlus, AffectNet-7, and AffectNet-8 respectively. Specifically, we compare with ResNet18-based methods, which are pre-trained on the face recognition dataset MS-Celeb-1M [67]: SAN [20], DMUE [29], RUL [28], DACL [55], MA-Net [56], DDA [26], and EAC [59]; with ResNet50-based methods: SCAN [27]. In addition, we also compare with transformer-based approaches such as MVT [57], VTFF [58], and TransFER [60]. Finally, we also compare with PSR [25] using VGG-16. We list all the comparison results in table I. We achieve the highest performance on all datasets as shown in table I. From the results, our method outperforms all current SOTA methods for FER, including ResNet-based and transformer-based methods. Specifically, we exceed the baseline [44] by 3.69%, 2.93%, 2.5%, 2.78% accuracy on RAF-DB, FERPlus, AffectNet-7, AffectNet-8.
IV-D1 Comparison with SL methods
@@
Considering the latest SL methods, we surpass EAC [59] by 2.06% and 1.79% on RAF-DB and AffectNet-7 respectively, PSR [25] by 1.36% on FERPlus, and DMUE [29] by 1.28% on AffectNet-8. These experimental results show the effectiveness of our proposed method in the FER task and demonstrate that converting ambiguous annotation results into one-hot labels using fully-supervised learning is not ideal. Compared to PLL, these CNN-based SL methods are misled by the ambiguous labels in the label space, which hinders the performance of the model. Furthermore, the transformer-based SL method suffers from the same problems. Although MVT [57] employs ViT as the backbone, it does not utilize unsupervised pre-training. Nevertheless, it remains susceptible to ambiguous labels in the label space, resulting in potential impacts on feature learning quality, which is consistent with the hypothesis proposed in [68]. The same problem exists with VTFF [58] and TransFER [60].
IV-D2 Comparison with PLL methods
@@
We also compare with several representative PLL methods: RC [62], LW [38], PICO [42], and CRPLL [63]. For a fair comparison, we use their optimal experimental settings. The results in table I are our implementation based on their open-source codes. The general PLL method for label disambiguation, e.g. RC and LW, relies on the robustness of the features, and from the results, for the FER task, RC, LW, and CRPLL are misled by ambiguous labels in the label space, which affects disambiguation due to the lack of robust facial expression feature representation. In addition, it is worth noting the result of PICO demonstrates that contrastive-based partial label learning does not apply to the FER task. As mentioned before, the contrastive-based approach is sensitive to data augmentation. Unlike objects, facial expressions have high inter-class similarities, which consist of five fixed facial action units; whereas typical data augmentation usually works on the lighting and contrast variation of the image, which only leads to extra redundant information in the FER task, and, sometimes, even to performance degradation.
IV-E Ablation Study
In this part, we conducted extensive ablation experiments, including: the effectiveness of PLL section IV-E1, the effectiveness of Pre-training section IV-E2, the feature-level effects of pre-trained SL, SSL section IV-E3, the confidence-level performance section IV-E4, the decoder and key modules section IV-E5, the ablation for threshold of revision confidence section IV-E6, the influence of $K$ section IV-E7, and the ablation for temperature $\tau$ section IV-E8.

### v4

@@ 3.3. Label Disambiguation
The label disambiguation with the transformer decoder mainly includes the label embedding regularization module and confidence revision module. An input image $x$ is fed into the pre-trained encoder to get visual token $\mathcal{F}^{HW\times d}$, where $HW$ denotes the number of visual tokens and dddenotes the feature dimension of the hidden state. We initialize the query embedding $\mathcal{Q}^{k\times d}$, where $k$ denotes the number of classes. In the training stage, the transformer decoder updates $\mathcal{Q}^{k\times d}$ according to visual tokens, which reflects the correlation of category-related features. Inspired by previous work (carion2020end; liu2021query2label), we link the category-related features to the learnable query embedding $\mathcal{Q}^{k\times d}$ and predict the confidence sector $C_{i}$ of the sample by their correspondence. Specifically, $C_{i}$ is obtained by inputting the logits from the last layer of the transformer decoder to the sigmoid function. However, differently, (1) Our method is designed for PLL and focuses on disambiguation by query to finally find the unique correct label with the highest confidence for an instance. (2) Besides querying the classification logits, we also query the label confidence. (3) To mitigate the problem of imbalance distribution in the FER dataset, we design a special regularization for label embedding.
3.3.1. Label Embedding Regularization
@@
It is worth noting that in addition to the ambiguity annotation these FER datasets also prevalently exhibit imbalanced distribution. In the training stage, it makes the model more biased towards majority classes, resulting in an uneven distribution of the feature space, and degrading model performance by classifying minority classes as head classes. In our work, we draw inspiration from  (wang2020understanding) and aim to achieve a uniform distribution of the query embedding over a hypersphere to mitigate the negative effects of class imbalance. Specifically, we compute a uniform loss for query embedding to pull the distance between different classes on the hypersphere. Specifically, we compute the optimal positions for different classes of query embedding $\left\{t_{i}\right\}_{i=1}^{K}$ before training and optimize it using the following Eq.(3):
(3) $\mathcal{L}_{uniform}\left(\left\{t_{i}\right\}_{i=1}^{K}\right)=\frac{1}{K}\sum_{i=1}^{K}\log\sum_{j=1}^{K}e^{t_{i}^{T}\cdot t_{j}/\tau},$
@@
where $\tau>0$ is a temperature hyperparameter, $K$ denotes the number of classes in FER dataset. We explore the impact of $K$ in the section 4.5.
3.3.2. Revision Confidence
@@
As previously mentioned, the robust representation in the feature space enhances PLL’s ability to disambiguate in the label space. For samples with ambiguous labels in the dataset, we initially set the pseudo targets for candidate label set $Y$ with a uniform distribution, represented by $C_{i}=\frac{1}{|Y_{i}|}Y_{i}$. During the update process, we continuously revise the confidence score $C_{i}$ for each batch using the following Eq.(4):
(4) $\binom{id_{1}:v_{1}}{id_{2}:v_{2}}=Top_{k}[\varphi*C_{i-1}],k=2,$
@@
where $\varphi=sigmoid(\sigma_{logits})$ is the response of sigmoid layer, and $\sigma_{logits}$ is the output of the transformer decoder. After that, we compute the $top2$ metrics of the $(C_{i-1}*\varphi)$ for updating confidence vector $C_{i}$ as the Eq.(5):
@@
(5) $C_{i}=\left\{\begin{array}[]{lcr}C_{i-1}\;\;\quad\quad\quad\quad\quad\quad\Delta{v}\leq threshold\\ C_{i-1}[id_{1}]=1\quad\quad\quad else\end{array}\right$.,
@@
where $\Delta{v}=|v_{1}-v_{2}|$, the threshold $\in$ [0,1). When $\Delta{v}>$ threshold, $C_{i}$ is a one-hot vector, where only the 1-th entry is 1. We treat it as the level of uncertainty and represent it with the top2 rank label in the candidate label set. Algorithm 1 is the pseudo-code of our method, which would converge to a stable status along the training iterations.
@@
0:ModelfModel\;f: Pre-trained model $f$;
@@
queryembeddingquery\;embedding: query $\mathcal{Q}$ ;
@@
   calculate Uniform distribution for query by Eq.(3);
   revise confidence vector $C_{i}$ by Eq.(4) and Eq.(5);
   calculate PLL loss by Eq.(2)
   update Network parameter of $f$ by Eq.(2);
@@
4. Experiments
In this section, we first describe the FER datasets and experiment implementation details. To evaluate our proposed method, we compare our method with baseline and other related methods including fully-supervised learning: PSR (vo2020pyramid), DDA (farzaneh2020discriminant), SCN (wang2020suppressing), SCAN (gera2021landmark), DMUE (she2021dive), RUL (zhang2021relative), DACL (farzaneh2021facial), MA-Net (zhao2021learning), MVT (li2021mvt), VTFF (ma2021facial), EAC (zhang2022learn), TransFER (xue2021transfer); partial label learning: RC (feng2020provably), LW (wen2021leveraged), PICO (wang2022pico), CRPLL (wu2022revisiting).
4.1. Datasets
RAF-DB (li2019reliable) contains around 30K diverse facial images with single or compound expressions labeled by 40 trained annotators via crowdsourcing. In our experiments, we use 12271 images for training and 3068 images for testing.
FERPlus (BarsoumICMI2016) is an extended dataset of the standard emotion FER dataset with 28709 training, 3589 validation, and 3589 test images. It provides a set of new labels for each image labeled by 10 human annotators. According to previous work, we just use the training part and test part.
AffectNet (DBLP:journals/corr/abs-1708-03985) is one of the most challenging FER datasets with manually labeled 440k images. As per prior research, we divide it into AffectNet-7 and AffectNet-8, with the latter having one more expression of contempt and 3667 training and 500 test images compared to the former.
Table 1. Pre-training setting, Part I.
@@ 4.1. Datasets
Table 2. Pre-training setting, Part II.
@@
Table 3. Fine-tuning setting of Maskfeat on FER datasets.
@@
4.2. Baselines and Experiment Setup
We adopt ViT-B/16 (dosovitskiy2020image) pre-trained on Imagenet-1K as our baseline. For fairness, we keep the same parameters as our fine-tuning part in all settings, which will be described later. In terms of SL methods, we adopt their optimal settings to make fair comparisons. For data preprocessing, we keep all the images with the size of $224\times 224$ and use MobileFacenet (chen2018mobilefacenets) to obtain aligned face regions for RAF-DB and FERPlus. For AffectNet, alignment was obtained via the landmarks provided in the data. In the pre-training part, we use four Tesla V100 GPUs to pre-train the ViT for 600 epochs with batch size 192 and learning rate 2e-4. We set the mask ratio of Maskfeat to 0.4. In fine-tuning part, we use only two Tesla V100 GPUs. We utilize the AdamW optimizer, setting the batch size to 320 and the initial learning rate to 1e-4 with the weight decay of 5e-2. The pre-training and fine-tuning protocol that we use is based on (wei2022masked). In the PLL setup, for the FER dataset without crowdsourcing results, we construct the candidate set for it based on LW (wen2021leveraged).
4.2.1. Details for pre-training and fine-tuning
Pre-training and Fine-tuning. Table 2 and Table 2 summarizes the pre-training configurations of different pre-trained methods.  Table 3 shows the fine-tuning configurations on the FER dataset. We use three FER datasets to pre-train the encoder part and adopt 2 transformer blocks in the decoder part.
4.2.2. Time/Memory Consumption of Proposed Method
We trained 600 epochs using 4 46GB Tesla A40, requiring  1.5 days for pre-training, which takes longer and relatively more memory resources. The related pre-trained checkpoint will be released to help with FER downstream tasks. As for fine-tuning, we do an ablation on RAF-DB with batch size=72/128/256, and it takes around 1.43/1.2/1.18 hours and around 11/16/30 GB of memory, which is similar to DMUE(she2021dive))/MA-NET(zhao2021learning)/EAC(zhang2022learn). However, we use similar resources to achieve better performance.
4.3. Comparison with the SOTA FER Methods
@@
We conduct extensive comparisons with the SOTA FER methods on RAF-DB, FERPlus, AffectNet-7, and AffectNet-8, respectively. Specifically, we compared ResNet18-based method, which is pre-trained on face recognition dataset MS-Celeb-1M (guo2016ms), such as SAN (wang2020suppressing), DMUE (she2021dive), RUL (zhang2021relative), DACL (farzaneh2021facial), MA-Net (zhao2021learning), DDA (farzaneh2020discriminant), EAC (zhang2022learn). ResNet50-based: SCAN (gera2021landmark), DR-FER (10375788), and FSLDL (shin2024noisy). In addition, we also compare Transformer-based such as MVT (li2021mvt), VTFF (ma2021facial), TransFER (xue2021transfer), MAE-Face (ma2023unified), and TA-FER (10149390). And only PSR (vo2020pyramid) using VGG-16. We list all the comparison results in Table 4. We achieve the highest performance on all datasets as shown in Table 4. From the results, our method outperforms all current SOTA methods for FER, including ResNet-based and Transformer-based methods. Specifically, we exceed the baseline (dosovitskiy2020image) by 3.69%, 2.93%, 2.5%, 2.78% accuracy on RAF-DB, FERPlus, AffectNet-7, AffectNet-8.
Table 4. Comparisons with the state-of-the-art methos on the RAF-DB, FERPlus, AffectNet-7 and AffectNet-8, respectively. Accuracy (%) is reported on the test dataset. denotes our implementation. “PR” denotes pre-train.
Method Backbone RAF-DB FERPlus AffectNet-7 AffectNet-8
PSR (vo2020pyramid) VGG-16 88.98 89.75 63.77 60.68
DDA (farzaneh2020discriminant) ResNet-18 86.90 - 62.34 -
SCN (wang2020suppressing) ResNet-18 87.03 88.01 63.40 60.23
SCAN (gera2021landmark) ResNet-50 89.02 89.42 65.14 61.73
FDRL (Ruan_2021_CVPR) ResNet-18 89.42 - - -
DMUE (she2021dive) ResNet-18 88.76 88.64 - 62.84
RUL (zhang2021relative) ResNet-18 88.98 - 61.43 -
DACL (farzaneh2021facial) ResNet-18 87.78 88.39 65.20 60.75
MA-Net (zhao2021learning) ResNet-18 88.40 - 64.53 60.96
MVT (li2021mvt) DeIT-S/16 88.62 89.22 64.57 61.40
VTFF (ma2021facial) ResNet-18+ViT-B/32 88.14 88.81 64.80 61.85
TransFER (xue2021transfer) CNN+ViT 90.91 89.76 66.23
EAC (zhang2022learn) ResNet-18 89.99 89.64 65.32 -
DR-FER (10375788) ResNet-50 90.61 91.91 67.54 63.60
FSLDL (shin2024noisy) ResNet-50 90.97 - 65.60 -
TA-FER (10149390) ViT-B/32 89.12 90.67 65.17 -
MAE-Face (ma2023unified) ViT-B/32 - - 69.51 66.65
RC (feng2020provably) ResNet-18 88.92 88.53 64.87 61.63
LW (wen2021leveraged) ResNet-18 88.70 88.10 64.53 61.46
PICO (wang2022pico) ResNet-18 88.53 87.93 64.35 61.38
CRPLL (wu2022revisiting) ResNet-18 88.47 88.01 64.32 61.37
Baseline (dosovitskiy2020image) ViT-B/16 88.36 88.18 64.61 61.34
ours w/o PR ViT-B/16 90.91 89.76 65.23 62.45
[…]
4.3.1. Comparison with SL methods
@@
Compared with the latest SL methods, we surpass EAC (zhang2022learn) by 2.06% and 1.79% on RAF-DB and AffectNet-7 respectively, PSR (vo2020pyramid) by 1.36% on FERPlus, and DMUE (she2021dive) by 1.28% on AffectNet-8. These experimental results show the effectiveness of our proposed method in the FER task and demonstrate that converting ambiguous annotation results into one-hot labels using fully-supervised learning is not ideal. Compared to PLL, these CNN-based SL methods are misled by the incorrect labels in the label space, which hinders the performance of the model. Furthermore, the Transformer-based SL method suffers from the same problems. Although MVT (li2021mvt) employs ViT as the backbone, it does not utilize unsupervised pre-training. Nevertheless, it remains susceptible to incorrect labels in the label space, resulting in potential impacts on feature learning quality, which is consistent with the hypothesis proposed in (liu2021self). The same problem exists with VTFF (ma2021facial) and TransFER (xue2021transfer). Other SL-based approaches, including SCAN (gera2021landmark), DR-FER (10375788), and FSLDL (shin2024noisy), also fall short of our method across all datasets. For example, our approach outperforms SCAN by 3.03%, 1.69%, and 1.97% on RAF-DB, FERPlus, and AffectNet-8, respectively. Similarly, although DR-FER benefits from advanced representation learning and achieves strong performance (e.g., 90.61% on RAF-DB and 91.91% on FERPlus), our method still improves over it by 1.44% and 0.80%, respectively. FSLDL is also clearly surpassed by our method on RAF-DB (90.97% vs. 92.05%), demonstrating that even stronger SL pipelines cannot compensate for the information loss caused by one-hot label conversion. Furthermore, although Transformer-based SL approaches such as TA-FER (10149390) achieve competitive results, they still underperform our model across all comparable datasets.
4.3.2. Comparison with SSL methods
Notably, MAE-Face (ma2023unified) performs large-scale masked-autoencoding pre-training on millions of external facial images, which essentially places it at a foundation-model scale rather than within the standard FER training regime. This makes direct comparison with supervised FER methods inherently unfair, as its performance is largely attributed to the capacity and generalization benefits brought by such extensive pre-training.
4.3.3. Comparison with PLL methods
@@
We add several representative PLL methods: RC (feng2020provably), LW (wen2021leveraged), PICO (wang2022pico), CRPLL (wu2022revisiting). For a fair comparison, we use their optimal experimental settings. The results in Table 4 are our implementation based on their open-source codes. The general PLL method for label disambiguation, such as RC and LW relies on the robustness of the features, and from the results, for the FER task, RC, LW, and CRPLL are misled by incorrect labels in the label space, which affects disambiguation due to the lack of robust facial expression feature representation. In addition, it is worth noting the result of PICO demonstrates that contrastive-based partial label learning does not apply to the FER task. As mentioned before, the contrastive-based approach is sensitive to data augmentation. Unlike objects, facial expressions have high inter-class similarities, which consist of five fixed facial action units. Whereas typical data augmentation usually works on the lighting and contrast variation of the image, which only leads to extra redundant information in the FER task, sometimes, even performance degradation.
4.4. Ablation Study